As part of a new fitness program, Elsa alternates running and walking in a park nearby her house. The formula 3w +5r represent the distance she walked and ran in the park last week, where w represents the amount of time she spent walking and r represent the amount of time she spent running. Elsa walked for 3 hours and she ran for 5 hours. How many miles total did Elsa walk and run last week?
Answer by stanbon(75887)

As part of a new fitness program, Elsa alternates running and walking in a park nearby her house. The formula 3w +5r represent the distance she walked and ran in the park last week, where w represents the amount of time she spent walking and r represent the amount of time she spent running. Elsa walked for 3 hours and she ran for 5 hours. How many miles total did Elsa walk and run last week?
---------------------------
3w + 4r = 3*3+5*5 = 9 + 25 = 34 miles
==========================================
Cheers,
Stan H.